OpenStudy (mathhelp346):

Each side of a square is increased by 2 centimeters, producing a new square whose area is 30 square centimeters. Find the length of the sides of the original square width.

OpenStudy (mathstudent55):

Original square side: s Increased side of each square: s + 2 Area of increased square: (s + 2)(s + 2) Area of increased square: 30 So (s + 2)(s + 2) = 30 s^2 + 4s + 4 = 30 s^2 + 4s - 26 = 0 Can you continue now?

OpenStudy (mathhelp346):

Yes thank you so much